1 pump can empty a tank in 8hr. and 15 min. another pump can empty the same tank in 11 hr how long will it take to empty the same tank with both pumps running at the same time.

Answer by sachi(548)

1 pump can empty a tank in 8hr. and 15 min=8+15/60 hr=33/4 hrs
so in 1 hr it will empty 4/33 parts
another pump can empty the same tank in 11 hr
so in 1 hr it will empty 1/11 parts
with both pumps running at the same time
in 1 hr both can empty 4/33+1/11=7/33 parts
so the time taken is 33/7 hrs=4 hrs 43 min approx to empty the same tank with both pumps running at the same time.